OpenStudy (anonymous):

amanda is running for city council, she needs to get 2over3 of her votes from senior citizens and 21000 votes in all to win , how many votes does she needs from voters other than the senior citizen??

OpenStudy (anonymous):

Since 2/3rd of votes come from senior citizens, the rest (1/3rd) should come from other people. The total is 21000, so it's simply\[21000\div 3 or 21000 * (1/3)\]

OpenStudy (texaschic101):

1/3 from others besides senior citizens = 1/3(21,000) = 21,000/3 = 7000 2/3 from seniors citizens = 2/3(21,000) = 42,000/3 = 14,000 total : 7,000 + 14,000 = 21,000